1. The current dividend yield on Clayton's Metals common stock is 3.2 percent. The company just...

1.

The current dividend yield on Clayton's Metals common stock is 3.2 percent. The company just paid a $1.48 annual dividend and announced plans to pay $1.54 next year. The dividend growth rate is expected to remain constant at the current level. What is the required rate of return on this stock?

7.25 percent

7.82 percent

8.08 percent

8.75 percent

8.39 percent

2.

Which one of the following is computed by dividing next year's annual dividend by the current stock price?

total yield

dividend yield

yield to maturity

growth rate

capital gains yield

3.

The bonds issued by Stainless Tubs bear an 8 percent coupon, payable semiannually. The bonds mature in 11 years and have a $1,000 face value. Currently, the bonds sell for $952. What is the yield to maturity?

7.92 percent

9.20 percent

8.69 percent

7.87 percent

8.08 percent

4.

Which one of the following applies to a premium bond?

Coupon rate > yield-to-maturity > current yield.
Coupon rate > current yield > yield to maturity.
Coupon rate < yield to maturity < current yield.

Yield to maturity > current yield > coupon rate.

Coupon rate = current yield = yield-to-maturity.

Please answer all the questions. Thank you.
Answer #1

1.

The current dividend yield is 3.2%, the dividend for next year is $1.54, as per dividend yield formula:

Dividend yield = Annual dividend / Current Market price

3.2%=$1.54/Price

Price = $1.54/3.2%

Price = $48.125

The growth rate is the growth of dividend from yr 1 to yr which is:

1.54-1.48 / 1.48

=0.04054

As per the dividend discount model,

P = D1 / r-g

48.125 = 1.54 / r-0.04054

=7.25%

The answer is a. 7.25%

2.

The answer is b.dividend yield as seen in the above part.

3.

The YTM is calculated using the Rate function in Ms-Excel:

=RATE(11,80,-952,1000)

=8.69%

The answer is c.8.69%

4.

The Answer is b. Coupon rate > current yield > yield to maturity.

In case of a premium bond, its coupon rate is higher than current yield and YTM.